OpenStudy (anonymous):

3(x+y)/4xy^2 when x equals 2/3 and y equals -2

OpenStudy (anonymous):

sub in x=2/3 and y=-2 into the equation... 3[(2/3)+(-2)]/4(2/3)(-2)^2 = 3[(2/3-2)]/[(8/3)(4)] (+ and - become negative) ( - squared is +) = 3(-4/3)/(32/3) (combine all terms in brackets) = (-4)/(32/3) (multiply terms in numerator to make it into one term) = 3/8